It is desired to filter a cell broth at a rate of 1700 liters/h on a rotary vacuum filter at a vacuum pressure of 70kPa. The cycle time for the drum will be 60 s, and the cake formation time (filtering time) will be 15 s. The broth to be filtered has a viscosity of 0.08 P and a cake solids (dry basis) per volume of filtrate of 10 g/l. From laboratory tests, the specific cake resistance has been determined to be 9 x 1010 cm/g. Determine the area of the filter that is required. The resistance of the filter can be neglected. 

 It is then desired to wash the product (antibiotic) out of the cake so that only 5% of the antibiotic in the cake is left after washing. The washing efficiency is 50%. Estimate the washing time per cycle that would be required. Assume that the filtered cake contains 70% water


1
Expert's answer
2020-07-13T19:06:28-0400

t/(V/A)="(\mu_0 \alpha\rho_c/2\Delta p)*(V/A)"

"A^2=" "\mu_0 \alpha\rho_c V^2/2 \Delta pt"

In applying this equation, it is helpful to focus on the area of the drum, which is where the cake is being formed and where filtrate is being obtained. Thus, A is the area of that part of the drum. We can calculate the volume of filtrate that needs to be collected during the cake formation time of 15 s:

V=1700*15*(1/3600)=7.083

"A^2=" (0.08*9*1010*10*7.0832*(1/104)*(1/102)4*(103))/(2*70*15)=1.72 m4

A=1.3115 m2

The area A’ of the entire rotary vacuum filter can be calculated from the cake formation time (15s) and the total cycle time (60s) as

A'=(60/15)*1.3115=5.246


 washing efficiency of a filter cake

R'=(1-E/100)n

where R’ is the weight fraction of solute remaining in the cake after washing (on the basis of R’ = 1.0 before washing), E is the percentage wash efficiency, and n is the volume of wash liquid per volume of liquid in the unwashed cake. Substituting R’ = 0.05 and E = 50% into this equation gives n=4.32

 the relationship between the washing time tw, and the cake formation time tf is given by tw=2nftf

where f is the ratio of volume Vr of residual liquid in the cake to the volume of filtrate Vf after time tf . Thus, we need to estimate the volume of residual liquid in the filter cake to determine tw . At the end of the 15 s cake formation time,  weight of dry cake solids filtered=10*7.083=70.83

weight of water=(70/30) *70.83=165.27 g

f=Vr/Vf=0.165/7.33=0.0225

tw=2nftf=2*4.32*0.0225*15=2.916 s